A sphere of radius 10.0cm is moulded into a uniform cylindrical wire of radius r calculate length of the wire
1
Expert's answer
2020-02-05T12:16:53-0500

When the sphere is moulded into a uniform cylindrical wire volume of the material remains constant

Volume of the sphere ="\dfrac{4}{3}\pi r'^3"

Volume of cylinder = "\pi r^2h"

"\to \pi r^2h= \dfrac{4}{3}\pi (r')^3"

"\to r^2h=(4/3)*1000"

"\to h= (4/3)*\dfrac{1000}{r^2}" length of the wire